The normality of 4.9% (w/w) H2SO4 solution having density 1.02g/ml is

Grade:

5 Answers

Arun
4 years ago
4.9 % acid by mass means 4.9g acid in 100 g solution. 
 Density = 1.02 g/mL
Volume of  solution = Mass /density
 = 100/1.02
 = 98.03 ml
Molar mass sulphuric acid H₂So₄ = 2*1 + 32 + 4*16 = 98g/mol
Number of moles of sulphuric acid = mass /  molar mass
=4.9 / 98
=0.05
Molarity = number of moles / volume of solution in litres
= 0.05 / 0.9803
= 0.051 M 
Normality = Molarity * Basicity
= 0.051*2 =0.102 N
Hence, the normality of 4.9% of H₂So₄ of density 1.02g/ml is 0.102N
 
 
Vikas TU
4 years ago
4.9% w/w H2SO4 means
4.9g H2SO4 in 100ml solution 
find molarity
m = (4.9 × 1000) / (98×100)
m= 0.5
then, find normality by multiplying molarity by valency factor.
valency factor for H2SO4 is 2
Nisha Sangwan
2 years ago
4.9 % acid by mass means 4.9g acid in 100 g solution. 
 Density = 1.02 g/mL
Volume of solution = Mass /density
 = 100/1.02
 = 98.03 ml
Molar mass sulphuric acid H₂So₄ = 2*1 + 32 + 4*16 = 98g/mol
Number of moles of sulphuric acid = mass / molar mass
=4.9 / 98
=0.05
Molarity = number of moles / volume of solution in litres
= 0.05 / 0.9803
= 0.051 M 
Normality = Molarity * Basicity
= 0.051*2 =0.102 N
Hence, the normality of 4.9% of H₂So₄ of density 1.02g/ml is 0.102N
Aniket Raj
2 years ago
4.9 % acid by mass means 4.9g acid in 100 g solution. 
 Density = 1.02 g/mL
Volume of solution = Mass /density
 = 100/1.02
 = 98.03 ml
Molar mass sulphuric acid H₂So₄ = 2*1 + 32 + 4*16 = 98g/mol
Number of moles of sulphuric acid = mass / molar mass
=4.9 / 98
=0.05
Molarity = number of moles / volume of solution in litres
= 0.05 / 0.09803
= 0.51 M 
Normality = Molarity * Basicity
= 0.51*2 =1.02 N
Hence, the normality of 4.9% of H₂So₄ of density 1.02g/ml is 1.02N
 
 
irshad hussain
2 years ago

Molarity={(w/w%)×10×(density in g/ml)}÷(Molar mass of solute in g)

 

M = (98*10*1.8)/98

 

= 18M

 

Normality = Molarity * nfactor(For H2SO4 2,Dibasic acid)

 

= 18*2

 

= 36N

 

Molality =(1000*Molarity)/{(1000*density in g/ml)-(Molar mass of solute in g*Molarity)}

 

=(1000*18)/(1000*1.8)-(98*18)

 

= 18000/36=500m